# Multi-Order Wavelet Derivative Transform for Deep Time Series Forecasting

### v1

@@ B.3 Implementation Details
All experiments are implemented in PyTorch [36] and executed on a single NVIDIA RTX A6000 GPU (40 GB). Optimization is carried out with Adam [20]; the initial learning rate $\eta$ is sampled log-uniformly from $[10^{-5},10^{-2}]$, the loss defined in Section 3.2.3 is minimized for training with early stop mechanism, and the batch size is tuned over $\{16,32,64,128\}$. Following FITS almost entirely, we use optuna [1] to carefully tune three hyper-parameters—look-back window length from $[192,1440]$, Wavelet decomposition scale $s\!\in\!\{1,2,3,4,5\}$, and Wavelet derivative order $k\!\in\!\{1,2,3,4,5\}$—choosing the best $L$ on the validation set to prevent information leakage, because longer contexts (e.g., $L\!=\!336$) typically improve most models whereas iTransformer [26] and FreTS [56] are comparatively insensitive. After fixing the optimal hyper-parameters, we retrain the model and report its test performance, while all baselines are re-run with their officially recommended configurations to ensure a fair comparison.
@@
where $m$ is the data periodicity. OWA is a specialized metric employed in the M4 competition [28], which facilitates the comparison of SMAPE and MASE against a Naïve2 baseline.
Appendix C Additional Results
C.1 Long-term Forecasting
Table 6 presents the long-term forecasting performance of our proposed WaveTS model compared to various contemporary methodologies across multiple datasets and forecasting horizons (96, 192, 336, 720). The reported results are averaged over 5 runs with different random seeds (see standard deviations in Table 7). Two-tailed paired Student’s t-tests ($\mathbf{\alpha=0.05}$) applied to the per-run metrics show that WaveTS’s improvements compared with previous baselines are statistically significant for most datasets and forecasting horizons (all $\mathbf{p<0.05}$). Specifically, WaveTS consistently outperforms existing FRL-based forecasters, such as FITS [49] and DeRiTS [7] achieving the best MSE and MAE in most datasets. For instance, in the Electricity dataset, WaveTS secures the lowest average MSE of 0.160 and MAE of 0.253 across all forecasting horizons. Across all datasets, WaveTS maintains the lowest overall average MSE of 0.314 and MAE of 0.338, demonstrating a substantial improvement of approximately 5.1% in MSE compared to the next best model, FITS [49]. Additionally, WaveTS achieves top performance in the Traffic and Exchange datasets, further highlighting its robust ability to model complex temporal patterns.
@@ C.1 Long-term Forecasting
WaveTS FITS DeRiTS WPMixer TexFilter TimeMixer iTransformer PatchTST FEDformer
@@
ETTm1 96 0.301 0.344 0.306 0.348 0.691 0.541 0.309 0.346 0.321 0.361 0.375 0.400 0.334 0.368 0.329 0.367 0.379 0.419
192 0.338 0.365 0.340 0.369 0.708 0.550 0.350 0.369 0.367 0.387 0.360 0.380 0.377 0.391 0.367 0.385 0.426 0.441
336 0.367 0.384 0.373 0.388 0.719 0.558 0.372 0.394 0.401 0.409 0.391 0.397 0.426 0.420 0.399 0.410 0.445 0.459
720 0.416 0.412 0.424 0.419 0.742 0.572 0.430 0.422 0.477 0.448 0.454 0.441 0.491 0.459 0.454 0.439 0.543 0.490
Avg 0.356 0.376 0.361 0.381 0.715 0.555 0.365 0.383 0.391 0.401 0.395 0.405 0.407 0.410 0.387 0.400 0.448 0.452
ETTm2 96 0.162 0.252 0.165 0.256 0.227 0.308 0.170 0.254 0.175 0.258 0.176 0.258 0.180 0.264 0.175 0.259 0.203 0.287
192 0.215 0.292 0.219 0.294 0.284 0.338 0.228 0.293 0.240 0.301 0.238 0.300 0.250 0.309 0.241 0.302 0.269 0.328
336 0.263 0.326 0.271 0.328 0.339 0.370 0.290 0.330 0.311 0.347 0.300 0.341 0.311 0.348 0.305 0.343 0.325 0.366
720 0.335 0.373 0.352 0.382 0.434 0.419 0.367 0.390 0.414 0.405 0.390 0.397 0.412 0.407 0.402 0.400 0.421 0.415
Avg 0.244 0.311 0.252 0.315 0.321 0.359 0.264 0.317 0.285 0.328 0.276 0.324 0.288 0.332 0.281 0.326 0.304 0.349
ETTh1 96 0.367 0.391 0.374 0.396 0.625 0.531 0.368 0.394 0.382 0.402 0.376 0.401 0.386 0.405 0.374 0.399 0.376 0.420
192 0.404 0.414 0.407 0.416 0.665 0.550 0.419 0.419 0.430 0.429 0.430 0.420 0.441 0.436 0.417 0.422 0.420 0.448
336 0.427 0.432 0.430 0.436 0.710 0.574 0.438 0.433 0.472 0.451 0.480 0.459 0.487 0.458 0.431 0.436 0.459 0.465
720 0.440 0.455 0.435 0.458 0.730 0.608 0.446 0.460 0.481 0.473 0.497 0.480 0.503 0.491 0.445 0.463 0.506 0.507
Avg 0.410 0.423 0.412 0.427 0.682 0.566 0.418 0.427 0.441 0.439 0.446 0.440 0.454 0.448 0.417 0.435 0.440 0.460
ETTh2 96 0.267 0.333 0.273 0.339 0.380 0.400 0.281 0.336 0.293 0.343 0.289 0.340 0.297 0.349 0.302 0.348 0.358 0.397
192 0.332 0.375 0.334 0.377 0.442 0.435 0.350 0.380 0.374 0.396 0.371 0.390 0.380 0.400 0.388 0.400 0.429 0.439
336 0.349 0.396 0.356 0.398 0.465 0.461 0.374 0.405 0.417 0.430 0.385 0.414 0.428 0.432 0.426 0.433 0.496 0.487
720 0.380 0.428 0.384 0.427 0.452 0.459 0.412 0.432 0.449 0.460 0.413 0.434 0.427 0.445 0.431 0.446 0.463 0.474
Avg 0.332 0.383 0.337 0.385 0.435 0.439 0.354 0.388 0.383 0.407 0.365 0.395 0.383 0.407 0.387 0.407 0.436 0.449
Electricity 96 0.131 0.227 0.145 0.242 0.275 0.362 0.148 0.240 0.147 0.245 0.153 0.247 0.148 0.240 0.181 0.270 0.193 0.308
192 0.146 0.240 0.157 0.252 0.277 0.364 0.161 0.250 0.160 0.251 0.167 0.256 0.162 0.253 0.188 0.274 0.201 0.315
336 0.162 0.256 0.174 0.269 0.291 0.376 0.177 0.265 0.173 0.267 0.185 0.277 0.178 0.269 0.204 0.293 0.214 0.329
720 0.200 0.288 0.213 0.301 0.329 0.402 0.215 0.302 0.210 0.309 0.226 0.311 0.225 0.317 0.246 0.324 0.246 0.355
Avg 0.160 0.253 0.172 0.266 0.293 0.376 0.175 0.264 0.172 0.268 0.183 0.273 0.178 0.270 0.205 0.290 0.213 0.327
Exchange 96 0.086 0.204 0.109 0.235 0.143 0.271 0.102 0.220 0.091 0.211 0.090 0.201 0.088 0.208 0.089 0.205 0.148 0.278
192 0.177 0.300 0.229 0.350 0.240 0.355 0.202 0.310 0.186 0.305 0.189 0.304 0.178 0.298 0.179 0.299 0.271 0.315
336 0.322 0.411 0.400 0.463 0.387 0.456 0.360 0.433 0.380 0.449 0.350 0.443 0.331 0.417 0.366 0.435 0.460 0.427
720 0.860 0.693 1.095 0.781 0.940 0.938 1.041 0.923 0.896 0.712 0.929 0.720 0.852 0.698 0.901 0.714 1.195 0.695
Avg 0.361 0.402 0.458 0.457 0.427 0.505 0.426 0.471 0.388 0.421 0.390 0.417 0.362 0.405 0.383 0.413 0.518 0.429
Traffic 96 0.382 0.266 0.401 0.280 0.961 0.542 0.431 0.312 0.430 0.294 0.461 0.286 0.395 0.268 0.462 0.295 0.587 0.366
192 0.394 0.270 0.415 0.286 0.973 0.547 0.411 0.310 0.452 0.307 0.473 0.299 0.416 0.276 0.466 0.296 0.604 0.373
336 0.409 0.278 0.429 0.290 0.959 0.536 0.443 0.311 0.470 0.316 0.500 0.306 0.430 0.282 0.482 0.304 0.621 0.383
720 0.447 0.298 0.468 0.308 1.010 0.556 0.505 0.329 0.498 0.323 0.501 0.312 0.466 0.305 0.514 0.322 0.626 0.382
Avg 0.408 0.278 0.428 0.291 0.976 0.545 0.448 0.316 0.462 0.310 0.484 0.301 0.427 0.283 0.481 0.304 0.609 0.376
Weather 96 0.167 0.223 0.145 0.199 0.216 0.270 0.168 0.205 0.162 0.207 0.169 0.212 0.174 0.214 0.177 0.218 0.217 0.296
192 0.210 0.258 0.190 0.243 0.264 0.304 0.209 0.270 0.210 0.250 0.210 0.255 0.221 0.254 0.225 0.259 0.276 0.336
336 0.256 0.294 0.238 0.282 0.312 0.335 0.263 0.289 0.265 0.290 0.257 0.290 0.278 0.296 0.278 0.297 0.339 0.380
720 0.315 0.336 0.310 0.332 0.380 0.375 0.339 0.339 0.342 0.340 0.340 0.343 0.358 0.347 0.354 0.348 0.403 0.428
Avg 0.237 0.278 0.230 0.266 0.293 0.321 0.235 0.283 0.245 0.272 0.244 0.275 0.258 0.278 0.259 0.280 0.309 0.360
Overall Avg 0.314 0.338 0.331 0.348 0.518 0.458 0.336 0.356 0.346 0.356 0.345 0.354 0.344 0.354 0.350 0.357 0.410 0.400
@@
Figure 7: Comparison of MSE reduction of several recent SOTA deep time series forecasting models.
Furthermore, it is worth emphasizing that our 5.1% reduction in MSE represents a significant leap forward. In many recently proposed forecasting methods [53, 41, 13, 46, 27, 22], improvements over baselines tend to remain below the 5%, as illustrated in Figure 7, indicating relatively marginal performance gains. Notably, the values shown in Figure 7 reflect each model’s average MSE improvement (or lack thereof) over its strongest baseline across a variety of datasets and forecasting horizons. In fact, some methods even fail to outperform the baseline when averaged over all experimental settings. Surpassing this margin underscores that WaveTS’s improvement is not marginal, highlighting its capacity in long-term forecasting.
C.2 Short-term Forecasting
@@
C.3 The Effectiveness of Backcasting while Forecasting
Table 10 compares each model trained only on the forecast window (F) with the same model trained to reconstruct the past and forecast the future (BF). MS is MambaSimple [10] for short). For every frequency-representation learner—WaveTS, FITS [49] and DeRiTS [7]—adding the backcast term consistently lowers both MSE and MAE at all four horizons on every dataset. The same trend holds for the linear state-space model MambaSimple [10], whose frequency-domain stem is a diagonal (element-wise) linear filter; details and the proof that such SSMs are FRL-equivalent are given in Appendix A.2. The gain comes from an architectural match: FRL modules apply invertible, element-wise linear transforms in the frequency domain, so the transformed spectrum still contains a recoverable copy of the original coefficients. Training with BF forces the network to exploit that embedded history, providing a richer supervisory signal, stabilizing optimization, and ultimately producing more accurate forecasts.
@@ C.3 The Effectiveness of Backcasting while Forecasting
C.4 Selection of the Branch Aggregation Strategy
@@
As illustrated in Figure 2, our proposed WaveTS is constructed upon a multi-branch architecture, wherein each branch corresponds to a specific order of Wavelet derivation. Upon obtaining the learned representations from each branch, the fusion and utilization of these representations presents a nontrivial issue. Consequently, this paper introduces two branch aggregation strategies aimed at identifying the optimal solution based on numerical results. The results, summarized in Table 11, indicate that concatenation along the temporal dimension consistently outperforms concatenation along the channel dimension. These findings demonstrate that the multi-scale patterns excavated by each branch are time-aware, reflecting temporal variations while aligning with the channel independence of the WDT within WaveTS. A similar phenomenon is also observed in experiments on the similar multi-branch DeRiTS [7].
C.5 Hyperparameter Sensitivity Analysis on Additional Datasets
We evaluate the sensitivity of our model to two key hyperparameters, the order $O$ and the scale $S$, on additional datasets: ETTm1, Weather, and ECL. Table 12 reports forecasting performance when varying $O$ (with $S=5$ fixed), and Table 13 shows the results for different $S$ (with $O=2$ fixed). We also provide results on even larger $O$ and $S$ in Table 8. Despite altering these hyperparameters, the changes in forecasting accuracy are relatively minor, indicating that our model maintains robust performance under varying settings.

### v2

@@ C.3 Implementation Details
All experiments are implemented in PyTorch (Paszke et al., 2019) and executed on a single NVIDIA RTX A6000 GPU (40 GB). Optimization is carried out with Adam (Kingma and Ba, 2014); the initial learning rate $\eta$ is sampled log‑uniformly from $10^{-5},10^{-2}]$, the loss defined in Section 3.2.3 is minimized for training with early stop mechanism, and the batch size is tuned over $\{16,32,64,128\}$. Following FITS almost entirely, we use optuna (Akiba et al., 2019) to carefully tune three hyper‑parameters—look‑back window length from $[192,1440]$, Wavelet decomposition scale $s\!\in\!\{1,2,3,4,5\}$, and Wavelet derivative order $k\!\in\!\{1,2,3,4,5\}$—choosing the best $L$ on the validation set to prevent information leakage, because longer contexts (e.g., L=336L\!=\!336) typically improve most models whereas iTransformer (Liu et al., 2024c) and FreTS (Yi et al., 2023b) are comparatively insensitive. After fixing the optimal hyper‑parameters, we retrain the model and report its test performance, while all baselines are re‑run with their officially recommended configurations to ensure a fair comparison.
@@
where $m$ is the data periodicity. OWA is a specialized metric employed in the M4 competition (Makridakis et al., 2018), which facilitates the comparison of SMAPE and MASE against a Naïve2 baseline.
Appendix D Additional Results
D.1 Long-term Forecasting
Table 6 presents the long-term forecasting performance of our proposed WaveTS model compared to various contemporary methodologies across multiple datasets and forecasting horizons (96, 192, 336, 720). The reported results are averaged over 5 runs with different random seeds (see standard deviations in Table 7). Two-tailed paired Student’s t-tests ($\mathbf{\alpha=0.05}$) applied to the per-run metrics show that WaveTS’s improvements compared with previous baselines are statistically significant for most datasets and forecasting horizons (all $\mathbf{p<0.05}$). Specifically, WaveTS consistently outperforms existing FRL-based forecasters, such as FITS (Xu et al., 2024) and DeRiTS (Fan et al., 2014) achieving the best MSE and MAE in most datasets. Across all datasets, WaveTS maintains the lowest overall average MSE of 0.314 and MAE of 0.338, demonstrating a substantial improvement of approximately 5.1% in MSE compared to the next best model, FITS (Xu et al., 2024). Additionally, WaveTS achieves top performance in the ETT datasets, further highlighting its robust ability to model complex temporal patterns.
@@ D.1 Long-term Forecasting
WaveTS (Ours) FITS (2024) DeRiTS (2014) WPMixer (2025) TexFilter (2024a) TimeMixer++ 2025b iTransformer (2024c) PatchTST (2023) DLinear (2023)
@@
ETTm1 96 0.301 0.344 0.306 0.348 0.691 0.541 0.309 0.346 0.321 0.361 0.313 0.354 0.334 0.368 0.329 0.367 0.346 0.374
192 0.338 0.365 0.340 0.369 0.708 0.550 0.350 0.369 0.367 0.387 0.356 0.380 0.377 0.391 0.367 0.385 0.382 0.391
336 0.367 0.384 0.373 0.388 0.719 0.558 0.372 0.394 0.401 0.409 0.386 0.402 0.426 0.420 0.399 0.410 0.415 0.415
720 0.416 0.412 0.424 0.419 0.742 0.572 0.430 0.422 0.477 0.448 0.452 0.437 0.491 0.459 0.454 0.439 0.473 0.451
Avg 0.356 0.376 0.361 0.381 0.715 0.555 0.365 0.383 0.391 0.401 0.377 0.393 0.407 0.410 0.387 0.400 0.404 0.408
ETTm2 96 0.162 0.252 0.165 0.256 0.227 0.308 0.170 0.254 0.175 0.258 0.170 0.245 0.180 0.264 0.175 0.259 0.193 0.293
192 0.215 0.292 0.219 0.294 0.284 0.338 0.228 0.293 0.240 0.301 0.229 0.291 0.250 0.309 0.241 0.302 0.284 0.361
336 0.263 0.326 0.271 0.328 0.339 0.370 0.290 0.330 0.311 0.347 0.303 0.343 0.311 0.348 0.305 0.343 0.382 0.429
720 0.335 0.373 0.352 0.382 0.434 0.419 0.367 0.390 0.414 0.405 0.373 0.399 0.412 0.407 0.402 0.400 0.558 0.525
Avg 0.244 0.311 0.252 0.315 0.321 0.359 0.264 0.317 0.285 0.328 0.269 0.320 0.288 0.332 0.281 0.326 0.354 0.402
ETTh1 96 0.367 0.391 0.374 0.396 0.625 0.531 0.368 0.394 0.382 0.402 0.361 0.403 0.386 0.405 0.374 0.399 0.397 0.412
192 0.404 0.414 0.407 0.416 0.665 0.550 0.419 0.419 0.430 0.429 0.416 0.441 0.441 0.436 0.417 0.422 0.446 0.441
336 0.427 0.432 0.430 0.436 0.710 0.574 0.438 0.433 0.472 0.451 0.430 0.434 0.487 0.458 0.431 0.436 0.489 0.467
720 0.440 0.455 0.435 0.458 0.730 0.608 0.446 0.460 0.481 0.473 0.467 0.451 0.503 0.491 0.445 0.463 0.513 0.510
Avg 0.410 0.423 0.412 0.427 0.682 0.566 0.418 0.427 0.441 0.439 0.419 0.432 0.454 0.448 0.417 0.435 0.461 0.457
ETTh2 96 0.267 0.333 0.273 0.339 0.380 0.400 0.281 0.336 0.293 0.343 0.276 0.328 0.297 0.349 0.302 0.348 0.340 0.394
192 0.332 0.375 0.334 0.377 0.442 0.435 0.350 0.380 0.374 0.396 0.342 0.379 0.380 0.400 0.388 0.400 0.482 0.479
336 0.349 0.396 0.356 0.398 0.465 0.461 0.374 0.405 0.417 0.430 0.346 0.398 0.428 0.432 0.426 0.433 0.591 0.541
720 0.380 0.428 0.384 0.427 0.452 0.459 0.412 0.432 0.449 0.460 0.392 0.415 0.427 0.445 0.431 0.446 0.839 0.661
Avg 0.332 0.383 0.337 0.385 0.435 0.439 0.354 0.388 0.383 0.407 0.339 0.380 0.383 0.407 0.387 0.407 0.563 0.519
Electricity 96 0.131 0.227 0.145 0.242 0.275 0.362 0.148 0.240 0.147 0.245 0.135 0.222 0.148 0.240 0.181 0.270 0.210 0.302
192 0.146 0.240 0.157 0.252 0.277 0.364 0.161 0.250 0.160 0.251 0.147 0.235 0.162 0.253 0.188 0.274 0.210 0.305
336 0.162 0.256 0.174 0.269 0.291 0.376 0.177 0.265 0.173 0.267 0.164 0.245 0.178 0.269 0.204 0.293 0.223 0.319
720 0.200 0.288 0.213 0.301 0.329 0.402 0.215 0.302 0.210 0.309 0.212 0.310 0.225 0.317 0.246 0.324 0.258 0.350
Avg 0.160 0.253 0.172 0.266 0.293 0.376 0.175 0.264 0.172 0.268 0.165 0.253 0.178 0.270 0.205 0.290 0.225 0.319
Exchange 96 0.086 0.204 0.109 0.235 0.143 0.271 0.102 0.220 0.091 0.211 0.085 0.214 0.088 0.208 0.089 0.205 0.088 0.218
192 0.177 0.300 0.229 0.350 0.240 0.355 0.202 0.310 0.186 0.305 0.175 0.313 0.178 0.298 0.179 0.299 0.176 0.315
336 0.322 0.411 0.400 0.463 0.387 0.456 0.360 0.433 0.380 0.449 0.316 0.420 0.331 0.417 0.366 0.435 0.313 0.427
720 0.860 0.693 1.095 0.781 0.940 0.938 1.041 0.923 0.896 0.712 0.851 0.689 0.852 0.698 0.901 0.714 0.839 0.695
Avg 0.361 0.402 0.458 0.457 0.427 0.505 0.426 0.471 0.388 0.421 0.357 0.391 0.362 0.405 0.383 0.413 0.354 0.414
Traffic 96 0.382 0.266 0.401 0.280 0.961 0.542 0.431 0.312 0.430 0.294 0.392 0.253 0.395 0.268 0.462 0.295 0.650 0.396
192 0.394 0.270 0.415 0.286 0.973 0.547 0.411 0.310 0.452 0.307 0.402 0.258 0.416 0.276 0.466 0.296 0.598 0.370
336 0.409 0.278 0.429 0.290 0.959 0.536 0.443 0.311 0.470 0.316 0.428 0.263 0.430 0.282 0.482 0.304 0.605 0.373
720 0.447 0.298 0.468 0.308 1.010 0.556 0.505 0.329 0.498 0.323 0.441 0.282 0.466 0.305 0.514 0.322 0.645 0.394
Avg 0.408 0.278 0.428 0.291 0.976 0.545 0.448 0.316 0.462 0.310 0.416 0.264 0.427 0.283 0.481 0.304 0.625 0.383
Weather 96 0.167 0.223 0.145 0.199 0.216 0.270 0.168 0.205 0.162 0.207 0.155 0.205 0.174 0.214 0.177 0.218 0.195 0.252
192 0.210 0.258 0.190 0.243 0.264 0.304 0.209 0.270 0.210 0.250 0.201 0.245 0.221 0.254 0.225 0.259 0.237 0.295
336 0.256 0.294 0.238 0.282 0.312 0.335 0.263 0.289 0.265 0.290 0.237 0.265 0.278 0.296 0.278 0.297 0.282 0.331
720 0.315 0.336 0.310 0.332 0.380 0.375 0.339 0.339 0.342 0.340 0.312 0.334 0.358 0.347 0.354 0.348 0.345 0.382
Avg 0.237 0.278 0.230 0.266 0.293 0.321 0.235 0.283 0.245 0.272 0.226 0.262 0.258 0.278 0.259 0.280 0.265 0.315
Overall Avg 0.314 0.338 0.331 0.348 0.518 0.458 0.336 0.356 0.346 0.356 0.320 0.335 0.344 0.354 0.350 0.357 0.406 0.402
@@
Furthermore, it is worth emphasizing that our 5.1% reduction in MSE represents a significant leap forward. In many recently proposed forecasting methods (Yi et al., 2024a; Wang et al., 2024a; Hu et al., 2024a; Wang et al., 2024d; Liu et al., 2024d; Lin et al., 2024), improvements over baselines tend to remain below the 5%, indicating relatively marginal performance gains.
D.2 Short-term Forecasting
@@
D.3 The Effectiveness of Backcasting while Forecasting
Table 10 compares each model trained only on the forecast window (F) with the same model trained to reconstruct the past and forecast the future (BF). MS is MambaSimple (Gu and Dao, 2024) for short). For every frequency-representation learner: WaveTS, FITS (Xu et al., 2024) and DeRiTS (Fan et al., 2014), adding the backcast term consistently lowers both MSE and MAE at all four horizons on every dataset. The same trend holds for the linear state-space model MambaSimple (Gu and Dao, 2024), whose frequency-domain stem is a diagonal (element-wise) linear filter; details and the proof that such SSMs are FRL-equivalent are given in Appendix B.2. The gain comes from an architectural match: FRL modules apply invertible, element-wise linear transforms in the frequency domain, so the transformed spectrum still contains a recoverable copy of the original coefficients. Training with BF forces the network to exploit that embedded history, providing a richer supervisory signal, stabilizing optimization, and ultimately producing more accurate forecasts.
@@ D.3 The Effectiveness of Backcasting while Forecasting
D.4 Selection of the Branch Aggregation Strategy
@@
As illustrated in Figure 2, our proposed WaveTS is constructed upon a multi-branch architecture, wherein each branch corresponds to a specific order of Wavelet derivation. Upon obtaining the learned representations from each branch, the fusion and utilization of these representations presents a nontrivial issue. Consequently, this paper introduces two branch aggregation strategies aimed at identifying the optimal solution based on numerical results. The results, summarized in Table 11, indicate that concatenation along the temporal dimension consistently outperforms concatenation along the channel dimension. These findings demonstrate that the multi-scale patterns excavated by each branch are time-aware, reflecting temporal variations while aligning with the channel independence of the WDT within WaveTS. A similar phenomenon is also observed in experiments on the similar multi-branch DeRiTS (Fan et al., 2014).
D.5 Hyperparameter Sensitivity Analysis on Additional Datasets
We evaluate the sensitivity of our model to two key hyperparameters, the order $O$ and the scale $S$, on additional datasets: ETTm1, Weather, and ECL. Table 12 reports forecasting performance when varying $O$ (with $S=5$ fixed), and Table 13 shows the results for different $S$ (with $O=2$ fixed). We also provide results on even larger $O$ and $S$ in Figure 8. Despite altering these hyperparameters, the changes in forecasting accuracy are relatively minor, indicating that our model maintains robust performance under varying settings.